OpenStudy (roxy.girl):

help? 4t + 7 + 6t = -33

OpenStudy (qwertty123):

Okay so first you want to. Reorder the terms: 7 + 4t + 6t = -33 Combine like terms: 4t + 6t = 10t 7 + 10t = -33

OpenStudy (roxy.girl):

okay I knew that XD

OpenStudy (roxy.girl):

so add -7?

OpenStudy (qwertty123):

yes :) Add '-7' to each side of the equation. 7 + -7 + 10t = -33 + -7 Combine like terms: 7 + -7 = 0 0 + 10t = -33 + -7 10t = -33 + -7

OpenStudy (qwertty123):

Combine like terms: -33 + -7 = -40 10t = -40

OpenStudy (roxy.girl):

so divide by 10?

OpenStudy (qwertty123):

Yes Divide each side by '10'

OpenStudy (roxy.girl):

so -4

OpenStudy (qwertty123):

t = -4 yes :3

OpenStudy (roxy.girl):

thanks